OpenStudy (anonymous):

A triangle has side lengths 10,15 and 7. is the triangle acute, obtuse or right? explain

OpenStudy (raden):

just identify these : if 7^2 + 10^2 - 15^2 < 0 then it is obtuse if 7^2 + 10^2 - 15^2 > 0 then it is acute if 7^2 + 10^2 - 15^2 = 0 then it is right triangle

OpenStudy (raden):

whice one do you like ?

OpenStudy (raden):

which *

OpenStudy (anonymous):

i still dont get it @RadEn

OpenStudy (raden):

can you determine what is 7^2 + 10^2 - 15^2 ?

OpenStudy (anonymous):

49 100 225

OpenStudy (raden):

7^2 = 7 * 7 = 49 10^2 = 10 * 10 = 100 15^2 = 15 * 15 = 225 now, what is the value of 7^2 + 10^2 - 15^2 ?

OpenStudy (anonymous):

76

OpenStudy (raden):

7^2 + 10^2 - 15^2 = 49 + 100 - 225 = ...

OpenStudy (anonymous):

76?

OpenStudy (raden):

49 + 100 = 149 then subtracted by 225 149 - 225 = ... ?